# Oxygen Isotope Ratios in Hydrogen-Deficient Carbon Stars: A Correlation with Effective Temperature and Implications for White Dwarf Merger Outcomes

### v1

@@ 1 Introduction
Hydrogen-deficient Carbon (HdC) stars are a class of supergiant stars with unusual chemical compositions, characterized by an acute deficiency of hydrogen and an overabundance of carbon (Clayton, 1996, 2012; Lambert & Rao, 1994). Their peculiar chemical compositions suggest that they originate in the merger of a CO-core and an He-core white dwarf (Clayton et al., 2007) in what is known as the double-degenerate (DD) formation scenario. They are found in all old galactic sub-structures, the thick disk, bulge and halo, but also in the younger thin disk (Tisserand et al., 2024). Adding to their intrigue, HdC stars seem to comprise two varieties with similar chemical compositions but distinct photometric properties: R Coronae Borealis (RCB)-type stars that show erratic brightness variations resulting from dust ejection episodes (Tisserand et al., 2020), and dustless Hydrogen-deficient Carbon (dLHdC) stars that do not show dramatic brightness variations or any signs of dust formation (Tisserand et al., 2022). Why RCB stars form dust while dLHdC stars do not despite having similar chemical compositions is still a mystery.
@@
Several studies have attempted to identify chemical differences between RCBs and dLHdC stars as a step towards solving their dusty mystery. Using medium resolution K-band spectra of these stars covering the CO molecular absorption bandheads, Clayton et al. (2007) first noted that RCB stars and dLHdC stars possibly differ in their oxygen isotope ratios, with dLHdCs having 16O/18O , while RCBs having 16O/18O . This was confirmed with high-resolution spectra by García-Hernández et al. (2009, 2010). However, these studies were based on small samples of five RCB stars and two dLHdC stars. Although there were over fifty RCB stars known at the time, only four dLHdC stars were known, of which two were too warm to exhibit CO lines. However, this has changed recently with the discovery of 27 new dLHdC stars (Tisserand et al., 2022). Additionally, a large number of RCB stars have been observed at near-infrared wavelengths (Karambelkar et al., 2021, 2024), enabling spectroscopic comparisons of larger samples of RCB and dLHdC stars.
@@
Using medium resolution NIR spectra for the expanded sample of RCBs and dLHdCs, (Karambelkar et al., 2022) showed that dLHdC stars generally tend to have lower values of 16O/18O than RCBs, and suggested that this difference could be attributed to different masses or mass-ratios of the merging white-dwarfs Crawford et al. (2024). However, the oxygen isotope ratios measured from the medium resolution spectra have large uncertainties due to excessive blending with molecular CN lines in this wavelength range. This lack of precision precludes any further insights, such as quantitative comparisons with theoretical models or identification of trends within the data.
@@
In this paper, we present high resolution (R75000) K-band spectra of six RCB and six dLHdC stars. These are the highest resolution NIR spectra of the largest sample of RCB and dLHdC stars to date. We use these spectra to derive the oxygen isotope ratios and other elemental abundances of these stars and examine differences between RCB and dLHdC stars. The paper is structured as follows – Section 2 describes the spectroscopic observations, Section 3 describes the abundance measurement method, 4 examines the differences and trends of the abundance measurements. We conclude with a summary of our results in Section 5.
@@
Name | Class | Date (UT) | S/N | Telluric Standard | Exposure Time (min) |
HD 137613 | dLHdC | 2022-09-14 | 104 | HIP79881 | 5.0 |
A223* | dLHdC | 2022-09-14 | 127 | HIP93667 | 57.5 |
B42* | dLHdC | 2022-09-14 | 32 | HIP93667 | 20.0 |
HD 182040 | dLHdC | 2022-09-22 | 108 | HIP95793 | 3.0 |
C38* | dLHdC | 2022-09-23 | 73 | HIP93667 | 100.0 |
B566* | dLHdC | 2022-09-30 | 78 | HIP91137 | 90.0 |
WISE J1820+ | RCB | 2022-09-24 | 41 | HIP93667 | 24.0 |
WISE J1942+ | RCB | 2022-09-22 | 98 | HIP98953 | 50.0 |
IRAS 1813+ | RCB | 2022-09-22 | 149 | HIP95793 | 50.0 |
ASAS-RCB-18 | RCB | 2022-09-23 | 79 | HIP95793 | 13.5 |
V CrA | RCB | 2022-09-23 | 136 | HIP93470 | 10.0 |
AO Her | RCB | 2022-09-24 | 88 | HIP91315 | 10.0 |
ASAS-RCB-21 | RCB | 2022-09-24 | 81 | HIP95793 | 27.0 |
NSV 11154 | RCB | 2022-09-24 | 87 | HIP91315 | 54.0 |
WISE J1818+ | RCB | 2022-09-30 | 68 | HIP93667 | 7.5 |
@@
The following abbreviations are used throughout the text: WISE J194218.38-203247.5: WISE J1942+; IRAS 1813.5-2419: IRAS 1813+; WISE J181836.38-181732.8: WISE J1818+; WISE J182010.96-193453.4: WISE J1820+. The last two stars are referred to as WISE-ToI-222 and WISE-ToI-223 in Karambelkar et al. (2021); Tisserand et al. (2020).
@@ 2 Observations
Our data consists of high-resolution ( or 4 km/s) K-band spectra of nine RCB and six dLHdC stars acquired with the iSHELL spectrograph (Rayner et al., 2022) on the 3.2 m NASA Infrared Telescope Facility (IRTF) at Mauna Kea Observatory, Hawaii. The observations were obtained over six nights between 2022 September 14 and 2022 October 3, with the 0.375" slit, using the K order sorter in the K3 mode which covers the wavelength range of m. This region was chosen to include multiple bandheads of 12C16O and 12C18O, as well as several lines of 12C14N. The full log of observations is listed in Table 1. We reduced the spectra using the idl software spextool (Cushing et al., 2004), and were calibrated and corrected for telluric absorption using the xtellcor package (Vacca et al., 2003). Table 1 also lists the median signal-to-noise ratios for the spectra measured at the continuum level in the 2.3 – 2.4 m region, which ranges from 30 to 150.
For the RCB stars in our sample, we determined photometric phases at the epochs of the IRTF observations using time-series photometric information aggregated from various surveys by the DREAMS monitoring portal111https://dreams.anu.edu.au/monitoring/. Of the nine RCB stars, IRAS 1813+ and WISE J1820+ were observed in a decline, where their V-band brightness were fainter by and mag than maximum luminosity respectively. Subsequently, the spectra for these two stars have not been analyzed as the photospheric lines are heavily diluted by newly ejected dust clouds. Additionally, although it was observed at maximum light, V CrA was found to have a heavily diluted and featureless spectrum, likely due to a combination of its high estimated effective temperature ( K) and large K-band excess from circumstellar dust. ASAS-RCB-18 and ASAS-RCB-21 were recovering from decline phases and were both roughly 0.8 mag fainter than maximum luminosity, but this does not appear to appreciably dilute the photospheric spectrum, as we see lines that are stronger than or comparable to other stars. The remaining four RCB stars were observed at maximum light, bringing the total number of RCB stars analyzed in this work to six.
@@ 3.1 Notable Spectral Features
Our spectra cover several molecular absorption features such as the CN Red System of 12C14N, the first seven overtone bandheads of 12C16O, the first five overtones of 12C18O (listed in Table 3), and C2 absorption lines from the Phillips and Ballik-Ramsay systems. There is considerable blending of lines especially beyond 2.29 m. Figure 2 shows a region with prominent CN lines for both dLHdC and RCB stars, along with a CN synthetic spectrum generated for one star in each class. Figure 2 shows a region dominated by 12C16O and 12C18O lines for the dLHdC and RCB stars. We also have several atomic lines in our spectra, however most of them are severely blended with molecular features. A few prominent ones that we use for determining elemental abundances are listed in Table 3.
Our wavelength range also includes bandheads of 13C16O, but these are not detected in any star in our sample. This indicates a high 12C/13C ratio that is consistent with the DD scenario and other observations of HdC stars (Clayton et al., 2007; García-Hernández et al., 2009; Jeffery et al., 2011). We also searched for 1 – 0 lines of HF (using ExoMol linelist by Tennyson et al. (2024)) which were recently detected in the K-band spectrum of DY Per (García-Hernández et al., 2023) but found none. This is consistent with García-Hernández et al. (2023) observing a lack of these lines in the RCB WX CrA and other HdC stars. They note that the absence of HF lines is likely indicative of a severe H-deficiency rather than a F-deficiency in these stars, as the abundances of both affect the formation of HF in the atmosphere.
@@ 3.3 Spectral Modeling
We generated synthetic spectra using TSFitPy (Storm & Bergemann, 2023) – a python wrapper around the radiative transfer package Turbospectrum v20.0 (Alvarez & Plez, 1998; Plez, 2012; Gerber et al., 2023). We used the CO line list provided by B. Plez (priv. commun. and described in Goorvitch (1994)) and the ExoMol linelists for CN and C2 (Yurchenko et al., 2018; Tennyson et al., 2024). Atomic line lists are generated using the VALD database (Ryabchikova et al., 2015).
@@ 3.4 Spectral Fitting
We fit the models to our observed spectra to derive the abundances of different elements. In addition to the elemental abundances, the model fitting requires four additional stellar parameters – effective temperature (T), surface gravity (), microturbulence (), and macroturbulence (). The micro- and macroturbulence velocities are free parameters in our fitting and are estimated using the observed spectra (see below for details). We fix for all stars, computed by assuming M⊙, T K, and R⊙. We use a combination of an empirical color-temperature calibration (Crawford et al., 2023) along with estimates from spectral energy distribution fits in literature where available to estimate T. Table 2 lists the stellar parameters adopted for the stars analyzed here. We discuss how potential uncertainties in these values might propagate to derived abundances in Section 3.5.3.
@@
Finally, the oxygen abundance and isotope ratios are measured using a set of 12C16O and 12C18O lines. To measure the oxygen-isotope ratios, we first measure the goodness-of-fit () by fitting synthetic spectra to all selected absorption lines simultaneously for a range of input 16O/18O values, with the total oxygen abundance as a free parameter333This approach is used because TSFitPy does not natively support isotope ratio fitting., and and set to the values determined in the previous step. We then fit a 4th order polynomial to the vs 16O/18O curve to determine the best-fit 16O/18O value as the one with the minimum , and uncertainties on this using the criterion (see Fig. 4d). We then fix the 16O/18O ratio to its best-fit value, and measure the total oxygen abundance by fitting the CO absorption lines individually, similar to the measurement of C and N abundances (see Fig. 4c). We report the mean and the standard-deviation of the measured line-by-line total oxygen abundances.
@@
This process yields the abundances of nitrogen, oxygen, and the oxygen isotope ratio for each star in our sample, along with uncertainties for each quantity. Additionally, the atomic lines listed in Table 3 are used to obtain the abundances of Fe, Mg, Na, Ca, S, and Si. However, we are unable to measure accurate uncertainties for these elements because of the small number of their unblended absorption lines in our spectra. Based on the uncertainties for the CNO elements, we conservatively estimate an uncertainty of dex on the abundances of these elements.
Star | A(C) | A(N) | A(O) | [Fe] | [Mg] | [Na] | [Ca] | [S] | [Si] | ||
HD 137613 | 5500 | 9.86 0.12 | 9.24 0.06 | 8.72 0.08 | 0.52 | -0.2 | 0.4 | 0.3 | -0.2 | -0.2 | -0.4 |
HD 182040 | 5750 | 9.90 0.13 | 9.08 0.08 | 7.99 0.06 | 0.58 | -0.5 | 0.2 | -0.0 | -0.4 | -0.2 | -0.2 |
B566 | 5750 | 10.03 0.25 | 9.09 0.10 | 8.55 0.10 | 0.78 | -0.3 | 0.0 | 0.0 | -0.5 | -0.1 | 0.3 |
C38 | 5750 | 10.32 0.16 | 9.47 0.08 | 8.99 0.10 | 0.73 | 0.2 | 0.7 | 0.6 | 0.0 | 0.2 | 0.2 |
A223 | 6250 | 9.48 0.24 | 8.73 0.09 | 8.20 0.09 | 0.60 | -0.8 | -0.1 | -0.1 | -0.8 | -0.4 | -0.7 |
B42 | 5500 | 10.17 0.14 | 9.18 0.14 | 8.81 0.1 | 0.32 | -0.3 | 0.3 | 0.3 | 0.1 | -0.1 | 0.0 |
NSV11154 | 5250 | 8.90 0.06 | 7.45 0.15 | 7.19 0.10 | 7.18 | -0.8 | -1.2 | -1.6 | -1.5 | -0.1 | -0.8 |
ASAS-RCB-21 | 5000 | 8.86 0.06 | 7.74 0.13 | 7.77 0.16 | 7.54 | -0.9 | -1.4 | -1.2 | -1.1 | -0.6 | -1.1 |
ASAS-RCB-18 | 5000 | 9.15 0.05 | 7.83 0.19 | 8.01 0.14 | 9.73 | -0.7 | -1.2 | -1.0 | -0.9 | -0.9 | -0.4 |
WISE J1818+ | 5000 | 8.93 0.06 | 7.60 0.12 | 7.55 0.10 | 7.47 | -0.7 | -1.3 | -1.3 | -0.9 | -0.2 | -1.0 |
AO Her | 4750 | 8.95 0.05 | 6.95 0.11 | 7.76 0.14 | 69 | -1.0 | -1.3 | -1.5 | -1.7 | -0.3 | -0.1 |
WISE J1942+ | 4500 | 8.72 0.06 | 6.59 0.13 | 7.42 0.10 | 93 | -0.9 | -1.2 | -1.8 | -1.6 | 0.0 | -0.6 |
Note: Fe, Mg, Na, Ca, S, and Si abundances are reported relative to solar values of A(Fe) = 7.5, A(Mg) = 7.6, A(Na) = 6.3, A(Ca) = 6.4, A(S) = 7.2, A(Si) = 7.6, taken from Magg et al. (2022).
@@
We obtain consistent isotope ratio measurements for the two stars HD 137613 and HD 182040 that were analyzed previously using high-resolution spectra by García-Hernández et al. (2009). The CNO abundances are slightly different because the previous study used a fixed model abundance of , while we fit this using C2 lines. Medium-resolution spectra of all six dLHdC stars and three RCB stars in our sample were analyzed by Karambelkar et al. (2021). The isotope ratios of three of the dLHdCs and one of the RCBs are within the ranges reported from medium-resolution spectra, while the remaining ones are outside the ranges but consistent to a factor of few. The reported nitrogen abundances also have uncertainties spanning 2.5 dex in most cases. This highlights the importance of using high-resolution spectra to accurately measure abundances and isotope ratios in the K-band due to extensive blending of molecular features.
@@
We also see that all the RCBs have additional broadening of lines in their spectra, with values ranging from 6 – 8 km s-1. This is not observed in the dLHdCs, and the difference is visually evident in Figures 2, 2. Our instrumental resolution of 4 km s-1 was critical in detecting this dichotomy, and this could point at the presence of a more dynamic and turbulent atmosphere in RCBs which could be related to dust formation in these stars
Finally, we observe that the dLHdCs seem to have a higher metallicity than the RCBs, with the mean [Fe] for the former and [Fe] for the latter. All the stars in our sample have sub-solar metallicities, with the exception of C38 which has [Fe] . Figure 6 compares the abundances of all elements to the iron abundances, with all values relative to solar abundances. A detailed analysis of the abundance trends observed will be presented in a future work.

### v2

@@ 1 Introduction
Hydrogen-deficient Carbon (HdC) stars are a class of supergiant stars with unusual chemical compositions, characterized by an acute deficiency of hydrogen and an overabundance of carbon (Clayton, 1996, 2012; Lambert & Rao, 1994). Their peculiar chemical compositions suggest that they originate in the merger of a CO-core and an He-core white dwarf (Clayton et al., 2007) in what is known as the double-degenerate (DD) formation scenario. They are found in all old galactic sub-structures, the thick disk, bulge and halo, but also in the younger thin disk (Tisserand et al., 2024). Adding to their intrigue, HdC stars seem to comprise two varieties with similar chemical compositions but distinct photometric properties: R Coronae Borealis (RCB)-type stars that show erratic brightness variations resulting from dust obscuration episodes (Tisserand et al., 2020), and dustless Hydrogen-deficient Carbon (dLHdC) stars that do not show dramatic brightness variations or any signs of dust formation (Tisserand et al., 2022). Why RCB stars form dust while dLHdC stars do not despite having similar chemical compositions is still a mystery.
@@
Several studies have attempted to identify chemical differences between RCBs and dLHdC stars as a step towards solving their dusty mystery. Using medium resolution K-band spectra of these stars covering the CO molecular absorption bandheads, Clayton et al. (2007) first noted that RCB stars and dLHdC stars possibly differ in their oxygen isotope ratios, with dLHdCs having 16O/18O and RCBs having 16O/18O . This was confirmed with high-resolution spectra by García-Hernández et al. (2009, 2010). However, these studies were based on small samples of five RCB stars and two dLHdC stars. Although there were over fifty RCB stars known at the time, only four dLHdC stars were known, of which two were too warm to exhibit CO lines. However, this has changed recently with the discovery of 27 new dLHdC stars (Tisserand et al., 2022). Additionally, a large number of RCB stars have been observed at near-infrared wavelengths (Karambelkar et al., 2021, 2024), enabling spectroscopic comparisons of larger samples of RCB and dLHdC stars.
@@
Using medium resolution NIR spectra for the expanded sample of RCBs and dLHdCs, Karambelkar et al. (2022) showed that dLHdC stars generally tend to have lower values of 16O/18O than RCBs, and suggested that this difference could be attributed to different masses or mass-ratios of the merging white-dwarfs (Crawford et al., 2024). However, the oxygen isotope ratios measured from the medium resolution spectra have large uncertainties due to excessive blending with molecular CN lines in this wavelength range. This lack of precision precludes any further insights, such as quantitative comparisons with theoretical models or identification of trends within the data.
@@
In this paper, we present high resolution (R75000) K-band spectra of six RCB and six dLHdC stars. These are the highest resolution NIR spectra of the largest sample of RCB and dLHdC stars to date. We use these spectra to derive the oxygen isotope ratios and other elemental abundances of these stars and examine differences between RCB and dLHdC stars. The paper is structured as follows – Section 2 describes the spectroscopic observations, Section 3 describes the abundance measurement method, and Section 4 examines the differences and trends of the abundance measurements. We conclude with a summary of our results in Section 5.
@@
Name | Class | Date (UT) | S/N | Telluric Standard | Total Exposure Time (min) | Number of Coadds |
HD 137613 | dLHdC | 2022-09-14 | 104 | HIP79881 | 5.0 | 5 |
A223* | dLHdC | 2022-09-14 | 127 | HIP93667 | 60.0 | 6 |
B42* | dLHdC | 2022-09-14 | 32 | HIP93667 | 20.0 | 2 |
HD 182040 | dLHdC | 2022-09-22 | 108 | HIP95793 | 3.0 | 6 |
C38* | dLHdC | 2022-09-23 | 73 | HIP93667 | 100.0 | 10 |
B566* | dLHdC | 2022-09-30 | 78 | HIP91137 | 90.0 | 9 |
WISE J1820+ | RCB | 2022-09-24 | 41 | HIP93667 | 24.0 | 12 |
WISE J1942+ | RCB | 2022-09-22 | 98 | HIP98953 | 50.0 | 5 |
IRAS 1813+ | RCB | 2022-09-22 | 149 | HIP95793 | 50.0 | 5 |
ASAS-RCB-18 | RCB | 2022-09-23 | 79 | HIP95793 | 13.5 | 9 |
V CrA | RCB | 2022-09-23 | 136 | HIP93470 | 10.0 | 5 |
AO Her | RCB | 2022-09-24 | 88 | HIP91315 | 10.0 | 10 |
ASAS-RCB-21 | RCB | 2022-09-24 | 81 | HIP95793 | 28.0 | 14 |
NSV 11154 | RCB | 2022-09-24 | 87 | HIP91315 | 54.0 | 5 |
WISE J1818+ | RCB | 2022-09-30 | 68 | HIP93667 | 7.5 | 5 |
@@
The following abbreviations are used throughout the text: WISE J194218.38-203247.5: WISE J1942+; IRAS 1813.5-2419: IRAS 1813+; WISE J181836.38-181732.8: WISE J1818+; WISE J182010.96-193453.4: WISE J1820+. The last two stars are referred to as WISE-ToI-222 and WISE-ToI-223 in Karambelkar et al. (2021) and Tisserand et al. (2020).
@@ 2 Observations
Our data consists of high-resolution ( or 4 km s-1) K-band spectra of nine RCB and six dLHdC stars acquired with the iSHELL spectrograph (Rayner et al., 2022) on the 3.2 m NASA Infrared Telescope Facility (IRTF) at Mauna Kea Observatory, Hawaii. The observations were obtained under program 2022B065 (PI: Karambelkar) over six nights between 2022 September 14 and 2022 October 3, with the 0.375" slit, using the K order sorter in the K3 mode which covers the wavelength range of m. This region was chosen to include multiple bandheads of 12C16O and 12C18O, as well as several lines of 12C14N. The full log of observations is listed in Table 1, along with the median continuum signal-to-noise ratios in the 2.3 – 2.4 m region. Spectra were reduced using the idl software spextool (Cushing et al., 2004), and were calibrated and corrected for telluric absorption using the xtellcor package (Vacca et al., 2003). We used the in-built spextool module xspextool for dark and flat calibration, wavelength calibration using arc-lamp observations and extraction of the 1D spectrum in each science exposure. Multiple exposures of the same target were combined using the xcombspec submodule. Once telluric corrected, the different echelle orders were merged using the submodule xmergeorders, that scales the orders to match flux levels of overlapping wavelengths in successive orders. Continuum-normalized spectra were obtained using a custom python algorithm that identifies the continuum level by identifying regions relatively free of strong molecular absorption and fits a fourth-order polynomial to them. The spectra were doppler corrected to the rest frame of each star using radial velocity measurements and manually correcting offsets of up to 20–30 km s-1 caused by pulsations that are common in these stars (Tisserand et al., 2024).
For the RCB stars in our sample, we determined photometric phases at the epochs of the IRTF observations using time-series photometric information aggregated from various surveys by the DREAMS monitoring portal111https://dreams.anu.edu.au/monitoring/. Of the nine RCB stars, IRAS 1813+ and WISE J1820+ were observed in a decline, where their V-band brightness were fainter by and mag than maximum luminosity respectively. Subsequently, the spectra for these two stars have not been analyzed as the photospheric lines are heavily diluted due to a decrease of the photospheric to circumstellar dust shell flux ratio in the K-band caused by newly ejected dust clouds. Additionally, although it was observed at maximum light, V CrA was found to have a featureless spectrum, likely due to a combination of its high estimated effective temperature ( K) and large K-band excess from circumstellar dust. ASAS-RCB-18 and ASAS-RCB-21 were recovering from decline phases and were both roughly 0.8 mag fainter than maximum luminosity, but this does not appear to appreciably dilute the photospheric spectrum, as we see lines that are stronger than or comparable to other stars despite a similar V-band deficit as IRAS 1813+. To understand this apparent discrepancy, we examined mid-infrared NEOWISE W1 and W2 light curves (Mainzer et al., 2014), that track variations of emission from the circumstellar dust shell (Feast et al., 1997). The NEOWISE data shows that IRAS 1813+ was at its maximum mid-IR brightness at the time of the iSHELL observation, while ASAS-RCB-18 and 21 were at minimum mid-IR brightnesses, suggesting that the extent of dilution observed in the K-band is substantially lower for ASAS-RCB-18 and 21. For this reason, we include these objects in our analysis. The remaining four RCB stars were observed at maximum light, bringing the total number of RCB stars analyzed in this work to six.
@@ 3.1 Notable Spectral Features
Our spectra cover several molecular absorption features such as the CN Red System of 12C14N, the first seven overtone bandheads of 12C16O, the first five overtones of 12C18O (listed in Table 2), and C2 absorption lines from the Phillips and Ballik-Ramsay systems. There is considerable blending of lines especially beyond 2.29 m. Figure 2 shows a region with prominent CN lines for both dLHdC and RCB stars, along with a CN synthetic spectrum generated for one star in each class. Figure 2 shows a region dominated by 12C16O and 12C18O lines for the dLHdC and RCB stars. We also have several atomic lines in our spectra, however most of them are severely blended with molecular features. A few prominent ones that we use for determining elemental abundances are listed in Table 2. C I lines are only detected in the dLHdC stars in our sample.
Our wavelength range also includes bandheads of 13C16O, but these are not detected in any star in our sample. This indicates a high 12C/13C ratio that is consistent with the DD scenario and other observations of HdC stars (Clayton et al., 2007; García-Hernández et al., 2009; Jeffery et al., 2011). We also searched for 1 – 0 lines of HF (using ExoMol linelist by Tennyson et al. 2024) which were recently detected in the K-band spectrum of DY Per (García-Hernández et al., 2023) but found none. This is consistent with García-Hernández et al. (2023) observing a lack of these lines in the RCB WX CrA and other HdC stars. They note that the absence of HF lines is likely indicative of a severe H-deficiency rather than a F-deficiency in these stars, as the abundances of both affect the formation of HF in the atmosphere.
@@ 3.3 Spectral Modeling
We generated synthetic spectra using TSFitPy (Storm & Bergemann, 2023) – a python wrapper around the radiative transfer package Turbospectrum v20.0 (Alvarez & Plez, 1998; Plez, 2012; Gerber et al., 2023). We used the CO line list provided by B. Plez (priv. commun. and described in Goorvitch 1994) and the ExoMol linelists for CN and C2 (Yurchenko et al., 2018; Tennyson et al., 2024). Atomic line lists are generated using the VALD database (Ryabchikova et al., 2015). We found several irregularities in the C I line lists available in the literature (as noted previously by García-Hernández et al. 2009), and used a subset of lines from the VALD-Kurucz database that are present in our data. These are reported in Table 6.
While Turbospectrum v20.0 includes support for Non-Local Thermodynamic Equilibrium (NLTE) spectrum synthesis, the requisite departure coefficients and atomic models are only available for a limited set of atoms (Ca, Fe, Mg, Na, O, Si at the time of writing). NLTE calculations for molecules are not supported at this time, and given that the main focus of our work is to recover abundances from molecular lines, we only use Local Thermodynamic Equilibrium (LTE) spectrum synthesis in our analysis.
@@ 3.4 Spectral Fitting
We fit the models to our observed spectra to derive the abundances of different elements. In addition to the elemental abundances, the model fitting requires four additional stellar parameters – effective temperature (T), surface gravity (), microturbulence (), and macroturbulence (). The micro- and macroturbulence velocities are free parameters in our fitting and are estimated using the observed spectra (see below for details). We fix for all stars, computed by assuming M⊙, T K, and R⊙. We use a combination of an empirical color-temperature calibration (Crawford et al., 2023) along with estimates from spectral energy distribution fits in literature where available to estimate T – this is another fixed parameter that cannot be determined from our spectral fits. Table 3 lists the stellar parameters adopted for the stars analyzed here. We discuss how potential uncertainties in these values might propagate to derived abundances in Section 3.5.3.
@@
Finally, the oxygen abundance and isotope ratios are measured using a set of 12C16O and 12C18O lines. To measure the oxygen-isotope ratios, we first measure the goodness-of-fit () by fitting synthetic spectra to all selected absorption lines simultaneously for a range of input 16O/18O values, with the total oxygen abundance as a free parameter333This approach is used because TSFitPy does not natively support isotope ratio fitting., and and set to the values determined in the previous step. We then fit a 4th order polynomial to the vs 16O/18O curve to determine the best-fit 16O/18O value as the one with the minimum , and uncertainties on this using the criterion (see Fig. 4d). We then fix the 16O/18O ratio to its best-fit value, and measure the total oxygen abundance by fitting the CO absorption lines individually, similar to the measurement of C and N abundances. We report the mean and the standard-deviation of the measured line-by-line total oxygen abundances.
@@
This process yields the abundances of nitrogen, oxygen, and the oxygen isotope ratio for each star in our sample, along with uncertainties for each quantity. Additionally, the atomic lines listed in Table 2 are used to obtain the abundances of Fe, Mg, Na, Ca, S, and Si. C I lines in dLHdC stars are also analyzed as an additional estimate for the carbon abundances, discussed further in Section 3.5.4. We are unable to measure accurate uncertainties for these elements because of the small number of their unblended absorption lines in our spectra. Based on the uncertainties for the CNO elements, we conservatively estimate an uncertainty of dex on the abundances of these elements.
Star | A(C) | A(N) | A(O) | A(C)* | [Fe] | [Mg] | [Na] | [Ca] | [S] | [Si] | ||
HD 137613 | 5500 | 9.86 0.12 | 9.24 0.06 | 8.72 0.08 | 0.52 | 9.6 | -0.2 | 0.4 | 0.3 | -0.2 | -0.2 | -0.4 |
HD 182040 | 5750 | 9.90 0.13 | 9.08 0.08 | 7.99 0.06 | 0.58 | 9.7 | -0.5 | 0.2 | -0.0 | -0.4 | -0.2 | -0.2 |
B566 | 5750 | 10.03 0.25 | 9.09 0.10 | 8.55 0.10 | 0.78 | 9.9 | -0.3 | 0.0 | 0.0 | -0.5 | -0.1 | 0.3 |
C38 | 5750 | 10.32 0.16 | 9.47 0.08 | 8.99 0.10 | 0.73 | 10.2 | 0.2 | 0.7 | 0.6 | 0.0 | 0.2 | 0.2 |
A223 | 6250 | 9.48 0.24 | 8.73 0.09 | 8.20 0.09 | 0.60 | 9.3 | -0.8 | -0.1 | -0.1 | -0.8 | -0.4 | -0.7 |
B42 | 5500 | 10.17 0.14 | 9.18 0.14 | 8.81 0.1 | 0.32 | 9.8 | -0.3 | 0.3 | 0.3 | 0.1 | -0.1 | 0.0 |
NSV11154 | 5250 | 8.90 0.06 | 7.45 0.15 | 7.19 0.10 | 7.18 | – | -0.8 | -1.2 | -1.6 | -1.5 | -0.1 | -0.8 |
ASAS-RCB-21 | 5000 | 8.86 0.06 | 7.74 0.13 | 7.77 0.16 | 7.54 | – | -0.9 | -1.4 | -1.2 | -1.1 | -0.6 | -1.1 |
ASAS-RCB-18 | 5000 | 9.15 0.05 | 7.83 0.19 | 8.01 0.14 | 9.73 | – | -0.7 | -1.2 | -1.0 | -0.9 | -0.9 | -0.4 |
WISE J1818+ | 5000 | 8.93 0.06 | 7.60 0.12 | 7.55 0.10 | 7.47 | – | -0.7 | -1.3 | -1.3 | -0.9 | -0.2 | -1.0 |
AO Her | 4750 | 8.95 0.05 | 6.95 0.11 | 7.76 0.14 | 69 | – | -1.0 | -1.3 | -1.5 | -1.7 | -0.3 | -0.1 |
WISE J1942+ | 4500 | 8.72 0.06 | 6.59 0.13 | 7.42 0.10 | 93 | – | -0.9 | -1.2 | -1.8 | -1.6 | 0.0 | -0.6 |
Note: Fe, Mg, Na, Ca, S, and Si abundances are reported relative to solar values of A(Fe) = 7.5, A(Mg) = 7.6, A(Na) = 6.3, A(Ca) = 6.4, A(S) = 7.2, A(Si) = 7.6 (Magg et al., 2022). A(C)* represents the C abundance recovered from C I lines, only measurable for the dLHdC stars.
Note: In HdC stars, absolute abundances depend on the assumed C/He ratio, which cannot be directly measured from our spectra due to a lack of helium lines, potentially introducing uncertainties. The use of abundance ratios is more robust for analyzing the chemical composition of these stars (see Asplund et al. 2000).
@@
We obtain consistent isotope ratio measurements for the two stars HD 137613 and HD 182040 that were analyzed previously using high-resolution spectra by García-Hernández et al. (2009). The CNO abundances are slightly different because the previous study used a fixed model abundance of , while we fit this using C2 lines. Medium-resolution spectra of all six dLHdC stars and three RCB stars in our sample were analyzed by Karambelkar et al. (2021). The isotope ratios of three of the dLHdCs and one RCB are within the ranges reported from medium-resolution spectra, while the remaining do not match. The reported nitrogen abundances also have uncertainties spanning 2.5 dex in most cases. This highlights the importance of using high-resolution spectra to accurately measure abundances and isotope ratios in the K-band due to extensive blending of molecular features.
@@
We also see that all the RCBs have additional broadening of lines in their spectra, with values ranging from 6 – 8 km s-1. This is not observed in the dLHdCs, and the difference is visually evident in Figures 2 and 2. Our instrumental resolution of 4 km s-1 was critical in detecting this dichotomy, and this could point at the presence of a more dynamic and turbulent atmosphere in RCBs which could be related to dust formation in these stars.
Finally, we observe that the dLHdCs seem to have a higher metallicity than the RCBs, with the mean [Fe] for the former and [Fe] for the latter. All the stars in our sample have sub-solar metallicities, with the exception of C38 which has [Fe] .